# Distance Between Stochastic Linear Systems

### v1

@@ 3.5 Numerical Demonstrations
Figure 2: The chordal metric based type-1 Wasserstein distance $\hat{d}_{1}(P_{1},P_{2})$ between two systems $P_{1}(s)=\frac{1}{1+0.5s}$ and $P_{2}(s)=\frac{1}{(1+0.2s)(1+0.7s)}$ is shown here in blue colour. The upper bound using the support distance is given in red colour and its lower bound from Theorem˜3.18 is shown in magenta colour. Quantities that are functions of frequency is given by solid lines and their respective maximum values are shown in dashed horizontal lines in the same colour.
To demonstrate the proposed distance in the frequency domain, we consider two different systems with their nominal model transfer functions given by $P_{1}(s)=\frac{1}{1+0.5s}$ and $P_{2}(s)=\frac{1}{(1+0.2s)(1+0.7s)}$ respectively. A discretized frequency space $\Omega_{M}$ containing $M=1000$ points between $0.1,10^{3}]\textrm{ rad}/s$ was formed. At every frequency $\omega\in\Omega_{M}$, $N=100$ samples of frequency response data were generated by randomly perturbing the nominal frequency response at that frequency. Both the nominal frequency response and the empirical distribution containing the samples at every frequency were projected onto the Riemann sphere using the inverse Stereographic projection given by (9). The chordal distance metric based type-1 Wasserstein distance $\hat{d}_{1}(P_{1},P_{2})$ between two systems given by (32) was computed using the linear programming approach. The upper bound using the support distance was computed using Proposition˜3.16 and the lower bound was computed using Theorem˜3.18. The results are shown in Fig.˜2. The proposed frequency domain distance $\hat{d}_{1}(P_{1},P_{2})$ between the systems $P_{1}$ and $P_{2}$ was found to be $0.2916$. The corresponding upper and lower bounds computed using Proposition˜3.16 and Theorem˜3.18 were found to be $0.3075$ and $0.2831$ respectively.
@@ 3.6 Summary of Frequency Domain Distance
For SISO systems, the proposed distance measure in the frequency domain using the chordal distance measure mimicking the $\nu$-gap perspective comes in handy with nice pictorial illustration. We believe that this is just a starting point and there are several interesting future research extensions. Though in principle, we expect the theory to carry forward in a similar fashion from SISO to MIMO systems setting, we expect some inherent difficult that comes with higher dimensions to kick in. For example, we would be required to work with the Riemann sphere of higher dimensions and associated stereographic projections are more mathematically involved and complex in nature. Another problem is of dealing with the supremum with respect to the frequency parameter. This problem persists even in SISO and will continue to persist even in MIMO setting. To get around these issues, we would like to formulate and obtain an analogous distance measure between stochastic linear systems in the time domain setting using gap metric perspective where the process of taking supremum with respect to the frequency parameter would be absent. The exposition with the time domain setting will be carried out in the next section.
@@ 4 Problem Formulation in Time Domain
In the previous section, we elucidated how to formulate and compute distance between two stochastic systems modeled in the frequency domain. We now present an analogous approach to achieve the same objective in the time domain using the gap metric.
@@ 4.1 Uncertain Dynamical Systems
Consider two continuous time LTI dynamical systems living in the space of linear stochastic systems denoted by $\mathbf{\Sigma}$. Let the nominal models of both the systems be given in the following form:
(39) $\bar{\Sigma}_{i}:\Bigl{\{}\dot{x}_{i}(t)=A_{i}x_{i}(t)+B_{i}u_{i}(t),\quad y_{i}(t)=C_{i}x(t),\quad i\in\{1,2\},$
where for system $i\in\{1,2\}$ at time $t\in\mathbb{R}_{\geq 0}$, we refer to its system states as $x_{i}(t)\in\mathbb{R}^{n}$, the control inputs to the system as $u_{i}(t)\in\mathbb{R}^{m}$, the system outputs as $y_{i}(t)\in\mathbb{R}^{l}$ and the matrices $A_{i}\in\mathbb{R}^{n\times n}$, $B_{i}\in\mathbb{R}^{n\times m}$, $C_{i}\in\mathbb{R}^{l\times n}$. Real-world dynamical systems usually have some form of uncertainties associated with them either due to the lack of modelling tools or due to the inaccuracies of the modelling framework. Hence, in practice, all systems have inherent uncertainties affecting their evolution. We model the uncertainty affecting the evolution of the uncertain system $i\in\{1,2\}$ using the parameter $\theta_{i}\in\mathbb{R}^{p}$ with $p\leq(n^{2}+nm+ln)$ and $\theta$ directly affects the evolution of the perturbed system described as follows:
(40) $\Sigma_{i}(\theta_{i}):\left\{\begin{aligned} \dot{x}_{i}(t)&=A_{i}(\theta_{i})x_{i}(t)+B_{i}(\theta_{i})u_{i}(t),\\ y_{i}(t)&=C_{i}(\theta_{i})x_{i}(t).\end{aligned}\right$.
@@ 4.2 Gap Between Models
(41) $\mathrm{Gap}_{i}(\theta_{i}):=\delta_{g}(\bar{\Sigma}_{i},\Sigma_{i}(\theta_{i})).$
@@
(42) $\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}:=\delta_{g}(\bar{\Sigma}_{1},\bar{\Sigma}_{2}).$
@@ 4.3 Inferring Distribution of Perturbed Plant
(43) $\mathbf{f}_{i}=\left(\Phi_{i}\right)_{\#}\mathbf{f}_{\theta_{i}}=\left(\Phi_{i}\right)_{\#}\mathcal{N}(\mu_{\theta_{i}},\Sigma_{\theta_{i}}),$
@@
(44) $\left(\Phi_{i}\right)_{\#}\mathbf{f}_{\theta_{i}}(\mathbf{B})=\mathbb{P}(\Sigma_{i}(\theta_{i})\in\mathbf{B})=\mathbf{f}_{\theta_{i}}(\Phi^{-1}_{i}(\mathbf{B})).$
@@
(45a) $A(\theta)$ $=A^{0}+\sum_{k=1}^{d}\theta^{(k)}A^{(k)},$
(45b) $B(\theta)$ $=B^{0}+\sum_{k=1}^{d}\theta^{(k)}B^{(k)},$
(45c) $C(\theta)$ $=C^{0}+\sum_{k=1}^{d}\theta^{(k)}C^{(k)},$
@@
(46) $z\sim\mathcal{N}(J\mu_{\theta}+z^{0},\,J\Sigma_{\theta}J^{\top}),$
@@
$\mathrm{vec}(A(\theta))=\mathrm{vec}\left(A^{0}+\sum_{k=1}^{d}\theta^{(k)}A^{(k)}\right)=\mathrm{vec}(A^{0})+\sum_{k=1}^{d}\theta^{(k)}\mathrm{vec}(A^{(k)})$
$\mathrm{vec}(B(\theta))=\mathrm{vec}\left(B^{0}+\sum_{k=1}^{d}\theta^{(k)}B^{(k)}\right)=\mathrm{vec}(B^{0})+\sum_{k=1}^{d}\theta^{(k)}\mathrm{vec}(B^{(k)})$
$\mathrm{vec}(C(\theta))=\mathrm{vec}\left(C^{0}+\sum_{k=1}^{d}\theta^{(k)}C^{(k)}\right)=\mathrm{vec}(C^{0})+\sum_{k=1}^{d}\theta^{(k)}\mathrm{vec}(C^{(k)})$
@@
(47) $z=z^{0}+\sum_{k=1}^{d}\theta^{(k)}z^{(k)},\quad\text{where}\quad z^{(k)}:=\begin{bmatrix}\mathrm{vec}(A^{(k)})\\ \mathrm{vec}(B^{(k)})\\ \mathrm{vec}(C^{(k)})\end{bmatrix}$.
@@
With $\theta\sim\mathcal{N}(\mu_{\theta},\Sigma_{\theta})$ and since Gaussianity is preserved under affine transformations, we infer that $z\sim\mathcal{N}(J\mu_{\theta}+z^{0},\;J\Sigma_{\theta}J^{\top})$. Note that result also follows from (43).
@@ 4.4 Gap Metric Induced Type-$q$ Wasserstein Distance
(48a) $\mathrm{dist^{q}_{\Sigma_{1},\Sigma_{2},\delta_{g}}}$ $:=\underset{\pi_{\mathbf{f}}\in\Pi\left(\mathbf{f}_{1},\mathbf{f}_{2}\right)}{\inf}\mathbb{E}_{\pi_{\mathbf{f}}}\left[\delta_{g}(\Sigma_{1},\Sigma_{2})^{q}\right]$
(48b) $=\underset{\pi_{\mathbf{f}}\in\Pi\left(\mathbf{f}_{1},\mathbf{f}_{2}\right)}{\inf}\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\delta_{g}(\Sigma_{1},\Sigma_{2})^{q}\,\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2})$
(48c) $=\inf_{\pi_{\mathbf{P}}\in\Pi(\mathbf{f}_{P_{1}},\mathbf{f}_{P_{2}})}\int_{\mathrm{supp}\left(\mathbf{f}_{P_{1}}\right)\times\mathrm{supp}\left(\mathbf{f}_{P_{2}}\right)}\delta_{g}(P_{1},P_{2})^{q}\,\pi_{\mathbf{P}}(dP_{1},dP_{2}).$
@@
(49) $\pi_{\mathbf{P}}=(\mathbf{TF}\times\mathbf{TF})_{\#}\,\pi_{\mathbf{f}}.$
@@
(50) $\mathsf{dist}^{q}_{\Sigma_{1},\Sigma_{2},\delta_{g}}\leq\inf_{\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})}\left(\sup_{\begin{subarray}{c}\Sigma_{1}\in\mathrm{supp}(\mathbf{f}_{1})\\ \Sigma_{2}\in\mathrm{supp}(\mathbf{f}_{2})\end{subarray}}\delta_{g}(\Sigma_{1},\Sigma_{2})\right)^{q}.$
Proof 4.5.
From (48), we know that
$\mathsf{dist}^{q}_{\Sigma_{1},\Sigma_{2},\delta_{g}}=\inf_{\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})}\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\delta_{g}(\Sigma_{1},\Sigma_{2})^{q}\,\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2}).$
We see that for any $\Sigma_{1}\in\mathrm{supp}(\mathbf{f}_{1}),\Sigma_{2}\in\mathrm{supp}(\mathbf{f}_{2})$,
@@
$\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\delta_{g}(\Sigma_{1},\Sigma_{2})^{q}\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2})$ $\leq\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\bar{\delta}^{q}\,\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2})$
@@
Taking the infimum over all joint distributions $\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})$ gives (50).
@@
Let $\theta_{i}\sim\mathbf{f}_{\theta_{i}}$ for $i\in\{1,2\}$ and suppose that the distribution $\mathbf{f}_{i}$ of the perturbed model of system $i\in\{1,2\}$ be given by (43). Further, let assumption 4 hold true for both the systems. Assume that $\exists\mathbf{L}_{i}>0$ such that the gap metric satisfies
(51) $\delta_{g}\left(\Phi_{i}(\theta_{i}),\Phi_{i}(\theta^{\prime}_{i})\right)\leq\mathbf{L}_{i}\left\|\theta_{i}-\theta^{\prime}_{i}\right\|\quad\forall\theta_{i},\theta^{\prime}_{i}\sim\mathbf{f}_{\theta_{i}}.$
@@
(52) $\mathrm{dist^{q}_{\Sigma_{1},\Sigma_{2},\delta_{g}}}\leq\left(\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}+\sum^{2}_{i=1}\mathbf{L}_{i}\left(\mathbb{E}\left[\left\|\theta_{i}-\bar{\theta}_{i}\right\|^{q}\right]\right)^{\frac{1}{q}}\right)^{q}.$
Proof 4.7.
@@
(53) $\delta_{g}(\Phi_{1}(\theta_{1}),\Phi_{2}(\theta_{2}))$ $\leq\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}+\sum^{2}_{i=1}\mathbf{L}_{i}\left\|\theta_{i}-\bar{\theta}_{i}\right\|.$
For brevity of notations, let us define $Z:=\delta_{g}\left(\Phi_{1}(\theta_{1}),\Phi_{2}(\theta_{2})\right)$, and $U_{i}:=\left\|\theta_{i}-\bar{\theta}_{i}\right\|$. Let $\pi_{\theta}\in\Pi(\mathbf{f}_{\theta_{1}},\mathbf{f}_{\theta_{2}})$ be any joint distribution of $(\theta_{1},\theta_{2})$. Then, the corresponding joint distribution of the systems $\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})$ and $\pi_{\theta}\in\Pi(\mathbf{f}_{\theta_{1}},\mathbf{f}_{\theta_{2}})$ are related as
(54) $\pi_{\mathbf{f}}=(\Phi_{1}\times\Phi_{2})_{\#}\pi_{\theta}.$
@@
(55) $\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\delta_{g}(\Sigma_{1},\Sigma_{2})^{q}\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2})=\int_{\mathrm{supp}(\mathbf{f}_{\theta_{1}})\times\mathrm{supp}(\mathbf{f}_{\theta_{2}})}Z^{q}\,\pi_{\theta}(d\theta_{1},d\theta_{2}).$
@@
(56) $\left\|Z\right\|_{\mathcal{L}_{q}}\leq\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}+\sum^{2}_{i=1}\mathbf{L}_{i}\left\|U_{i}\right\|_{\mathcal{L}_{q}}.$
@@
(57) $\left\|U_{i}\right\|_{\mathcal{L}_{q}}=\left(\mathbb{E}_{\pi_{\theta}}[U_{i}^{q}]\right)^{1/q}=\left(\mathbb{E}_{\mathbf{f}_{\theta_{i}}}\left[\left\|\theta_{i}-\bar{\theta}_{i}\right\|^{q}\right]\right)^{1/q}.$
@@
Taking the infimum over all joint distributions $\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})$ and subsequently using (55) and the definition (48), we get
(58) $\left(\mathrm{dist^{q}_{\Sigma_{1},\Sigma_{2},\delta_{g}}}\right)^{\frac{1}{q}}\leq\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}+\sum^{2}_{i=1}\mathbf{L}_{i}\left(\mathbb{E}_{\mathbf{f}_{\theta_{i}}}\left[\left\|\theta_{i}-\bar{\theta}_{i}\right\|^{q}\right]\right)^{1/q}.$
Since $x\mapsto x^{q}$ is increasing on $\mathbb{R}_{\geq 0}$ for $q\geq 1$, we raise both sides of (58) to the power $q\geq 1$ to get (52) and the proof is complete.
In the following proposition, we will give a lower bound using the nominal gap distance and the deviation from nominal model. The reasoning will be very similar to that of Theorem˜3.11 in the frequency domain meaning that we have an analogous result in the time domain given by the following proposition.
@@
(59) $\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}^{q}$ $\geq\inf_{\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})}\left(\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}-\sum^{2}_{i=1}\mathbb{E}_{\pi_{\mathbf{f}}}[\mathrm{Gap}_{i}(\theta_{i})]\right)^{q}_{+}.$
Proof 4.9.

### v2

@@ 3.5 Numerical Demonstrations
Figure 2: The chordal metric based type-1 Wasserstein distance $\hat{d}_{1}(P_{1},P_{2})$ between two systems $P_{1}(s)=\frac{1}{1+0.5s}$ and $P_{2}(s)=\frac{1}{(1+0.2s)(1+0.7s)}$ is shown here in blue colour. The upper bound using the support distance is given in red colour and its lower bound from Theorem 3.6 is shown in magenta colour. Quantities that are functions of frequency is given by solid lines and their respective maximum values are shown in dashed horizontal lines in the same colour.
To demonstrate the proposed distance in the frequency domain, we consider two different systems with their nominal model transfer functions given by $P_{1}(s)=\frac{1}{1+0.5s}$ and $P_{2}(s)=\frac{1}{(1+0.2s)(1+0.7s)}$ respectively. A discretized frequency space $\Omega_{M}$ containing $M=1000$ points between $0.1,10^{3}]\textrm{ rad}/s$ was formed. At every frequency $\omega\in\Omega_{M}$, $N=100$ samples of frequency response data were generated by randomly perturbing the nominal frequency response at that frequency. Both the nominal frequency response and the empirical distribution containing the samples at every frequency were projected onto the Riemann sphere using the inverse Stereographic projection given by (9). The chordal distance metric based type-1 Wasserstein distance $\hat{d}_{1}(P_{1},P_{2})$ between two systems given by (32) was computed using the linear programming approach. The upper bound using the support distance was computed using Proposition 3.5 and the lower bound was computed using Theorem 3.6. The results are shown in Figure 2. The proposed frequency domain distance $\hat{d}_{1}(P_{1},P_{2})$ between the systems $P_{1}$ and $P_{2}$ was found to be $0.2916$. The corresponding upper and lower bounds computed using Proposition 3.5 and Theorem 3.6 were found to be $0.3075$ and $0.2831$ respectively.
@@ 3.6 Summary of Frequency Domain Distance
For SISO systems, the proposed distance measure in the frequency domain using the chordal distance measure mimicking the $\nu$-gap perspective comes in handy with nice pictorial illustration. We believe that this is just a starting point and there are several interesting future research extensions. Though in principle, we expect the theory to carry forward in a similar fashion from SISO to MIMO systems setting, we expect some inherent difficult that comes with higher dimensions to kick in. For example, we would be required to work with the Riemann sphere of higher dimensions and associated stereographic projections are more mathematically involved and complex in nature. Another problem is of dealing with the supremum with respect to the frequency parameter. This problem persists even in SISO and will continue to persist even in MIMO setting. To get around these issues, we would like to formulate and obtain an analogous distance measure between stochastic linear systems in the time domain setting using gap metric perspective where the process of taking supremum with respect to $\omega\in\Omega$ would be absent. The exposition with the time domain setting will be carried out in the next section.
@@ 4 Problem Formulation in Time Domain
In this section, we will present an analogous distance measure in the time domain using the gap metric.
@@ 4.1 Uncertain Dynamical Systems
Consider two continuous time LTI dynamical systems living in the space of linear stochastic systems denoted by $\mathbf{\Sigma}$. Let the nominal models of both systems $i\in\{1,2\}$ be given by
$\bar{\Sigma}_{i}:\Bigl{\{}\dot{x}_{i}(t)=A_{i}x_{i}(t)+B_{i}u_{i}(t),\quad y_{i}(t)=C_{i}x(t),$ (40)
where for system $i$ at time $t\in\mathbb{R}_{\geq 0}$, we refer to its system states as $x_{i}(t)\in\mathbb{R}^{n}$, the control inputs to the system as $u_{i}(t)\in\mathbb{R}^{m}$, the system outputs as $y_{i}(t)\in\mathbb{R}^{l}$ and the matrices $A_{i}\in\mathbb{R}^{n\times n}$, $B_{i}\in\mathbb{R}^{n\times m}$, $C_{i}\in\mathbb{R}^{l\times n}$. Real-world dynamical systems usually have some form of uncertainties associated with them either due to the lack of modelling tools or due to the inaccuracies of the modelling framework. Hence, in practice, all systems have inherent uncertainties affecting their evolution. We model the uncertainty affecting the evolution of the uncertain system $i\in\{1,2\}$ using the parameter $\theta_{i}\in\mathbb{R}^{p}$ with $p\leq(n^{2}+nm+ln)$ and $\theta$ directly affects the evolution of the perturbed system described as follows:
$\Sigma_{i}(\theta_{i}):\left\{\begin{aligned} \dot{x}_{i}(t)&=A_{i}(\theta_{i})x_{i}(t)+B_{i}(\theta_{i})u_{i}(t),\\ y_{i}(t)&=C_{i}(\theta_{i})x_{i}(t).\end{aligned}\right$. (41)
@@ 4.2 Gap Between Models
$\mathrm{Gap}_{i}(\theta_{i}):=\delta_{g}(\bar{\Sigma}_{i},\Sigma_{i}(\theta_{i})).$ (42)
@@
$\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}:=\delta_{g}(\bar{\Sigma}_{1},\bar{\Sigma}_{2}).$ (43)
@@ 4.3 Inferring Distribution of Perturbed Plant
$\mathbf{f}_{i}=\left(\Phi_{i}\right)_{\#}\mathbf{f}_{\theta_{i}}=\left(\Phi_{i}\right)_{\#}\mathcal{N}(\mu_{\theta_{i}},\Sigma_{\theta_{i}}),$ (44)
@@
$\left(\Phi_{i}\right)_{\#}\mathbf{f}_{\theta_{i}}(\mathbf{B})=\mathbb{P}(\Sigma_{i}(\theta_{i})\in\mathbf{B})=\mathbf{f}_{\theta_{i}}(\Phi^{-1}_{i}(\mathbf{B})).$ (45)
@@
$A(\theta)$ $=A^{0}+\sum_{k=1}^{d}\theta^{(k)}A^{(k)},$ (46a)
$B(\theta)$ $=B^{0}+\sum_{k=1}^{d}\theta^{(k)}B^{(k)},$ (46b)
$C(\theta)$ $=C^{0}+\sum_{k=1}^{d}\theta^{(k)}C^{(k)},$ (46c)
@@
$z\sim\mathcal{N}(J\mu_{\theta}+z^{0},\,J\Sigma_{\theta}J^{\top}),$ (47)
@@
$\mathrm{vec}(A(\theta))$ $=\mathrm{vec}\left(A^{0}+\sum_{k=1}^{d}\theta^{(k)}A^{(k)}\right)$
$=\mathrm{vec}(A^{0})+\sum_{k=1}^{d}\theta^{(k)}\mathrm{vec}(A^{(k)})$
$\mathrm{vec}(B(\theta))$ $=\mathrm{vec}\left(B^{0}+\sum_{k=1}^{d}\theta^{(k)}B^{(k)}\right)$
$=\mathrm{vec}(B^{0})+\sum_{k=1}^{d}\theta^{(k)}\mathrm{vec}(B^{(k)})$
$\mathrm{vec}(C(\theta))$ $=\mathrm{vec}\left(C^{0}+\sum_{k=1}^{d}\theta^{(k)}C^{(k)}\right)$
$=\mathrm{vec}(C^{0})+\sum_{k=1}^{d}\theta^{(k)}\mathrm{vec}(C^{(k)})$
@@
$z=z^{0}+\sum_{k=1}^{d}\theta^{(k)}z^{(k)},\quad\text{where}\quad z^{(k)}:=\begin{bmatrix}\mathrm{vec}(A^{(k)})\\ \mathrm{vec}(B^{(k)})\\ \mathrm{vec}(C^{(k)})\end{bmatrix}$. (48)
@@
With $\theta\sim\mathcal{N}(\mu_{\theta},\Sigma_{\theta})$ and since Gaussianity is preserved under affine transformations, we infer that $z\sim\mathcal{N}(J\mu_{\theta}+z^{0},\;J\Sigma_{\theta}J^{\top})$. Note that result also follows from (44). ∎
@@ 4.4 Gap Metric Induced Type-$q$ Wasserstein Distance
$\mathrm{dist^{q}_{\Sigma_{1},\Sigma_{2},\delta_{g}}}$
$:=\underset{\pi_{\mathbf{f}}\in\Pi\left(\mathbf{f}_{1},\mathbf{f}_{2}\right)}{\inf}\mathbb{E}_{\pi_{\mathbf{f}}}\left[\delta_{g}(\Sigma_{1},\Sigma_{2})^{q}\right]$ (49a)
$=\underset{\pi_{\mathbf{f}}\in\Pi\left(\mathbf{f}_{1},\mathbf{f}_{2}\right)}{\inf}\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\delta_{g}(\Sigma_{1},\Sigma_{2})^{q}\,\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2})$ (49b)
$=\inf_{\pi_{\mathbf{P}}\in\Pi(\mathbf{f}_{P_{1}},\mathbf{f}_{P_{2}})}\!\int_{\mathrm{supp}\left(\mathbf{f}_{P_{1}}\right)\times\mathrm{supp}\left(\mathbf{f}_{P_{2}}\right)}\!\!\!\delta_{g}(P_{1},P_{2})^{q}\,\pi_{\mathbf{P}}(dP_{1},dP_{2}).$ (49c)
@@
$\pi_{\mathbf{P}}=(\mathbf{TF}\times\mathbf{TF})_{\#}\,\pi_{\mathbf{f}}.$ (50)
@@
$\mathsf{dist}^{q}_{\Sigma_{1},\Sigma_{2},\delta_{g}}\leq\inf_{\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})}\left(\sup_{\begin{subarray}{c}\Sigma_{1}\in\mathrm{supp}(\mathbf{f}_{1})\\ \Sigma_{2}\in\mathrm{supp}(\mathbf{f}_{2})\end{subarray}}\delta_{g}(\Sigma_{1},\Sigma_{2})\right)^{q}.$ (51)
Proof.
For any $\Sigma_{1}\in\mathrm{supp}(\mathbf{f}_{1}),\Sigma_{2}\in\mathrm{supp}(\mathbf{f}_{2})$,
@@
$\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\delta_{g}(\Sigma_{1},\Sigma_{2})^{q}\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2})$
$\leq\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\bar{\delta}^{q}\,\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2})$
@@
Taking the infimum over all joint distributions $\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})$ gives (51). ∎
@@
Let $\theta_{i}\sim\mathbf{f}_{\theta_{i}}$ for $i\in\{1,2\}$ and suppose that the distribution $\mathbf{f}_{i}$ of the perturbed model of system $i\in\{1,2\}$ be given by (44). Further, let assumption 4.1 hold true for both the systems. Assume that $\exists\mathbf{L}_{i}>0$ such that the gap metric satisfies
$\delta_{g}\left(\Phi_{i}(\theta_{i}),\Phi_{i}(\theta^{\prime}_{i})\right)\leq\mathbf{L}_{i}\left\|\theta_{i}-\theta^{\prime}_{i}\right\|\quad\forall\theta_{i},\theta^{\prime}_{i}\sim\mathbf{f}_{\theta_{i}}.$ (52)
@@
$\mathrm{dist^{q}_{\Sigma_{1},\Sigma_{2},\delta_{g}}}\leq\left(\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}+\sum^{2}_{i=1}\mathbf{L}_{i}\left(\mathbb{E}\left[\left\|\theta_{i}-\bar{\theta}_{i}\right\|^{q}\right]\right)^{\frac{1}{q}}\right)^{q}.$ (53)
Proof.
@@
$\delta_{g}(\Phi_{1}(\theta_{1}),\Phi_{2}(\theta_{2}))$ $\leq\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}+\sum^{2}_{i=1}\mathbf{L}_{i}\left\|\theta_{i}-\bar{\theta}_{i}\right\|.$ (54)
Let us define $Z:=\delta_{g}\left(\Phi_{1}(\theta_{1}),\Phi_{2}(\theta_{2})\right)$, and $U_{i}:=\left\|\theta_{i}-\bar{\theta}_{i}\right\|$. Let $\pi_{\theta}\in\Pi(\mathbf{f}_{\theta_{1}},\mathbf{f}_{\theta_{2}})$ be any joint distribution of $(\theta_{1},\theta_{2})$. Then, the corresponding joint distribution of the systems $\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})$ and $\pi_{\theta}\in\Pi(\mathbf{f}_{\theta_{1}},\mathbf{f}_{\theta_{2}})$ are related as
$\pi_{\mathbf{f}}=(\Phi_{1}\times\Phi_{2})_{\#}\pi_{\theta}.$ (55)
@@
$\int_{\mathrm{supp}(\mathbf{f}_{1})\times\mathrm{supp}(\mathbf{f}_{2})}\delta_{g}(\Sigma_{1},\Sigma_{2})^{q}\pi_{\mathbf{f}}(d\Sigma_{1},d\Sigma_{2})$
$=$ $\int_{\mathrm{supp}(\mathbf{f}_{\theta_{1}})\times\mathrm{supp}(\mathbf{f}_{\theta_{2}})}Z^{q}\,\pi_{\theta}(d\theta_{1},d\theta_{2}).$ (56)
@@
$\left\|Z\right\|_{\mathcal{L}_{q}}\leq\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}+\sum^{2}_{i=1}\mathbf{L}_{i}\left\|U_{i}\right\|_{\mathcal{L}_{q}}.$ (57)
@@
$\left\|U_{i}\right\|_{\mathcal{L}_{q}}=\left(\mathbb{E}_{\pi_{\theta}}[U_{i}^{q}]\right)^{1/q}=\left(\mathbb{E}_{\mathbf{f}_{\theta_{i}}}\left[\left\|\theta_{i}-\bar{\theta}_{i}\right\|^{q}\right]\right)^{1/q}.$ (58)
@@
Taking the infimum over all joint distributions $\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})$ and subsequently using (4.4) and the definition (49), we get
$\left(\mathrm{dist^{q}_{\Sigma_{1},\Sigma_{2},\delta_{g}}}\right)^{\frac{1}{q}}\leq\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}+\sum^{2}_{i=1}\mathbf{L}_{i}\left(\mathbb{E}_{\mathbf{f}_{\theta_{i}}}\left[\left\|\theta_{i}-\bar{\theta}_{i}\right\|^{q}\right]\right)^{1/q}.$ (59)
Since $x\mapsto x^{q}$ is increasing on $\mathbb{R}_{\geq 0}$ for $q\geq 1$, we raise both sides of (59) to the power $q\geq 1$ to get (53) and the proof is complete. ∎
In the following proposition, we will give a lower bound using the nominal gap distance and the deviation from nominal model. The reasoning will be very similar to that of Theorem 3.3 in the frequency domain meaning that we have an analogous result in the time domain given by the following proposition.
@@
$\mathrm{dist}_{\Sigma_{1},\Sigma_{2},\delta_{g}}^{q}$ $\geq\inf_{\pi_{\mathbf{f}}\in\Pi(\mathbf{f}_{1},\mathbf{f}_{2})}\left(\mathrm{dist^{nom}_{\Sigma_{1},\Sigma_{2}}}-\sum^{2}_{i=1}\mathbb{E}_{\pi_{\mathbf{f}}}[\mathrm{Gap}_{i}(\theta_{i})]\right)^{q}_{+}.$ (60)
Proof.